OpenStudy (anonymous):

Each lap in the Indy 500 is 2.5 miles. A driver completes two laps in 1.5 minutes. The average speed during the second lap is 8 miles per hour (mph) faster than the average speed during the first lap. Find the driver's average speed for each of the two laps, to the nearest mile per hour. 204 mph; 212 mph 175 mph; 183 mph 183 mph; 191 mph 196 mph; 204 mph

OpenStudy (anonymous):

lapOne = lapTwo + 8 What are all the numbers at the bottom?

OpenStudy (anonymous):

Possible answers

OpenStudy (anonymous):

so then divide 1.5 by 2 and then divide 2.5/.75?

OpenStudy (anonymous):

no, that's if lapOne=lapTwo It's not a perfect split

OpenStudy (anonymous):

so then how exactly do i do it?

OpenStudy (anonymous):

1.5min = 5Miles So, 5 / 1.5 = x / 60 x = 200 they drive 200 mph on average, which is the average for the first two laps combined, because you know it's an 8mph difference the first lap is just 4 mph slower than the ave speed, and the second lap is 4 mph faster, final answers: 196 mph and 204 mph

OpenStudy (anonymous):

found that on yahoo answers, but it worked out fine.

OpenStudy (anonymous):

ohhh so its a proportion?

OpenStudy (anonymous):

Just to get the avg speed

OpenStudy (anonymous):

ok thanks